MLK Bank has an asset portfolio that consists of $150 million of 15-year, 7.5-percent-coupon, $1,000 bonds with annual coupon payments that sell at par.

b-1.

The duration of these bonds is 9.4892 years. What are the predicted bond prices in each of the four cases using the duration rule? (Do not round intermediate calculations. Round your answers to 2 decimal places. (e.g., 32.16))

              Bonds’ New Price
  At + 0.10% $     
  At − 0.10%     
  At + 2.0%     
  At − 2.0%     
b-2.

What is the amount of error between the duration prediction and the actual market values? (Negative amount should be indicated by a minus sign. Do not round intermediate calculations. Round your answers to 2 decimal places. (e.g., 32.16))

              Amount of Error
  At + 0.10% $     
  At − 0.10%     
  At + 2.0%     
  At − 2.0%     

PLEASE SHOW THE WORK ( I ALREADY POSTED THIS QUESTION AND THE ANSWER WAS WONG )
Answer #1

1
=(1-9.4892*0.1%)*1000=990.5108

2
=(1+9.4892*0.1%)*1000=1009.4892

3
=(1-9.4892*2%)*1000=810.216

4
=(1+9.4892*2%)*1000=1189.784

1
=PV(7.6%,15,-7.5%*1000,-1000)-990.5108=0.716635846638951

2
=PV(7.4%,15,-7.5%*1000,-1000)-1009.4892=-0.606998705041747

3
=PV(9.5%,15,-7.5%*1000,-1000)-810.216=33.220499947496

4
=PV(5.5%,15,-7.5%*1000,-1000)-1189.784=10.9676188694273